# Computing Distances on Graph Associahedra is Fixed-parameter Tractable

### v1

@@ 5 Further research
Another natural direction is to investigate whether Rotation Distance admits a polynomial kernel parameterized by $k$. So far, this is only known when the considered graph $G$ is a path, where even linear kernels are known[14, 30]; see Table 1. As an intermediate step, one may consider graphs of bounded degree, for which it seems plausible that 4.3 (restriction to few balls of bounded diameter) provides a helpful opening step.
[…]
Table 1: Known results and open problems about the (parameterized) complexity of the Rotation Distance problem, both on paths and general graphs.
Finally, Ito et al. [24] also proved the NP-hardness of a related problem called Combinatorial Shortest Path on Polymatroids, relying on the fact that graph associahedra can be realized as the base polytopes of some polymatroids [35]. To the best of our knowledge, the parameterized complexity of this problem has not been investigated.

### v2

@@ 6 Further research
We start with further research about computing distances on graph associahedra, and then on hypergraphic polytopes.
@@
Another natural direction is to investigate whether Rotation Distance admits a polynomial kernel parameterized by $k$. So far, this is only known when the considered graph $G$ is a path, where even linear kernels are known[17, 37]. As an intermediate step, one may consider graphs of bounded degree, for which it seems plausible that 4.3 (restriction to few balls of bounded diameter) provides a helpful opening step.
There are also several interesting questions about computing distances on hypergraphic polytopes. We proved in Theorem 5.2 that the problem is ${\sf W}[2]$-hard when parameterized by the target distance $\ell$. It makes sense to identify additional parameters that allow for FPT algorithms. A natural one is the rank of the hypergraph, that is, the maximum size of a hyperedge. Note that in the reduction of Theorem 5.2, the rank of the constructed hypergraph can be arbitrarily large. Note also that, in case the problem were FPT with the combined parameter distance and rank, it would not admit polynomial kernels unless ${\sf NP}\subseteq{\sf coNP}/{\sf poly}$ by our negative result in Theorem 5.11, since the rank of a hypergraph is never larger than its number of vertices. It is worth mentioning that the hardness result of Cardinal and Steiner on hypergraphic polytopes [13, Theorem 4.3] holds even if the input hypergraph has rank three (and additionally, bounded maximum degree).
Also, we proved in Theorem 5.8 that there exists some constant $c>0$ such that the Flip Distance Between Acyclic Orientations problem does not admit a polynomial-time $c\cdot\log(|V|+|{\mathcal{E}}|)$-approximation algorithm unless ${\sf P}={\sf NP}$, where $H=(V,{\mathcal{E}})$ is the input hypergraph. It remains open whether there exists a polynomial-time approximation algorithm matching (asymptotically) this ratio.